Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 14     Entry time: 01 Mar 2019, 14:18
Author: Adriano Violante 
Location: General 
Severity: Information 
Subject: Report after dispatch meeting 09.00 

SPB/SFX

  • Today Loop 8 could be ready (AE)
  • Agipd - visualization of data shows variable rep rate... correlations to be investigated.

FXE

SQS

SCS

MID

  • Fast diodes needed on the WE. HW and Software support required and ongoing.
  • CY called for a  meeting on monday to collect all the requirements for the epix det to work. The det will be then tested in the common lase room. Karabo interface change is required. It will take 3 weeks.
  • VAC interlock test in MID OPT to be performed. Marjian can support.

HED

  • Si(111) mono commissionig ongoing
  • VAC componennt in HED OPT being setup by KS
  • Digitizers for diodes being used.

OPT Lasers

XRO

​​XPD

​​VAC

AE

  • Timing update on monday in SASE2. BF sent an email to readiness with the details.

ITDM

DET

EET

CAS

  • Test Karabo 2.4.0. ongoing. List of PLCs AE==>CAS.
  • Problem with a solid attenuator valve in XTD1. Not enough air pressure to operate the valve properly. DFM(CAS) will contact XRO.

AoB

CAS and AE point out that there are many activities planned for the next shutdown and as soon as PSPO will issue the shutdown tasklist planning most probably priorities will need to be set by TT/ MGMT

ELOG V3.1.4-7c3fd00